OpenStudy (anonymous):

What are the solutions of the equation: z^2 - 12z + 36 = 0?

OpenStudy (linknissan):

http://www.1728.org/quadratc.htm

OpenStudy (linknissan):

put the values of indices in the calcuator

OpenStudy (anonymous):

Do NOT use the quadratic formula on this on. Much, much easier to factor this one. Always see if you can factor first.

OpenStudy (anonymous):

z^2 - 12z + 36 = 0 (z )(z ) = 0 Do you know how to figure out what goes in the ( ) ( ) 's ?

OpenStudy (anonymous):

nope

OpenStudy (linknissan):

u want to verify the answer or learn to solve it?

OpenStudy (anonymous):

z^2 - 12z + 36 = (z - 6)(z - 6) So we have (z - 6)(z - 6) = 0 z - 6 = 0 z = 6 There's only 1 solution on this one.

OpenStudy (anonymous):

z^2 - 12z + 36 = 0? Just factor it pretending the Z is an X so you get (X-6)(X-6) so x=6 or z=6 and since Z is only squared there will only be 2 solutions. ^^^^^ The answer before is right except there will be 2 solutions.

OpenStudy (anonymous):

i just want the and the explaintion lol... have no more patients anymore

OpenStudy (anonymous):

No there is only one solution. x = 6 What is the other solution?

OpenStudy (anonymous):

If you cannot understand it from Me and @BangkokGarrett then its hopeless. And yes sorry for this one there is only 1 solution sorry bangkok

OpenStudy (anonymous):

Factoring is tricky for a lot of students. Master that skill and you will benefit a great deal in future math courses.

OpenStudy (anonymous):

^^^ that's for sure. Factoring is one of the most common tools you will use throughout your math career.

OpenStudy (anonymous):

thank for your help i got it